OpenStudy (anonymous):

Solve the system by using the elimination method: 3x + 2y = 4 and -x + 3y = -5 ? A. (2, -1) B. (-2, 1) C. (-3, 2) D. none of the above

pooja195 (pooja195):

opposites cancel

pooja195 (pooja195):

3x + 2y = 4 --x + 3y = -5

pooja195 (pooja195):

What would u need to multiply by to cancel POSITIVE 3x

pooja195 (pooja195):

?(--x + 3y = -5 )

pooja195 (pooja195):

Any ideas?

pooja195 (pooja195):

OK illgive you that part 3(--x + 3y = -5 )

pooja195 (pooja195):

still waiting .-.

OpenStudy (anonymous):

you multiply each number by 3

pooja195 (pooja195):

EVERYTHING by 3

pooja195 (pooja195):

3(--x + 3y = -5 )